# PromptGAR: Flexible Promptive Group Activity Recognition

### v1

@@ Abstract
We present PromptGAR, a novel framework that addresses the limitations of current Group Activity Recognition (GAR) approaches by leveraging multi-modal prompts to achieve both input flexibility and high recognition accuracy. The existing approaches suffer from limited real-world applicability due to their reliance on full prompt annotations, the lack of long-term actor consistency, and under-exploration of multi-group scenarios. To bridge the gap, we proposed PromptGAR, which is the first GAR model to provide input flexibility across prompts, frames, and instances without the need for retraining. Specifically, we unify bounding boxes, skeletal keypoints, and areas as point prompts and employ a recognition decoder for cross-updating class and prompt tokens. To ensure long-term consistency for extended activity durations, we also introduce a relative instance attention mechanism that directly encodes instance IDs. Finally, PromptGAR explores the use of area prompts to enable the selective recognition of the particular group activity within videos that contain multiple concurrent groups. Comprehensive evaluations demonstrate that PromptGAR achieves competitive performances both on full prompts and diverse prompt inputs, establishing its effectiveness on input flexibility and generalization ability for real-world applications.
@@ 1 Introduction
Group Activity Recognition (GAR) [8] is fundamentally important for video and event understanding, and it is widely used in areas such as video analytics, human computer interaction, and security systems. GAR processes videos and annotated information including bounding boxes, skeletons, ball trajectories, and optical flows to determine the group activity label. Building on prior research [55, 32, 15, 29, 19, 50, 36, 38], we aim to not only enhance group activity recognition accuracy but also to create a more flexible architecture capable of handling diverse prompt, frame, and instance inputs.
@@
Figure 1: PromptGAR Architecture. A sequence of frames $I$ is processed by the video encoder, yielding RGB features $\textbf{F}_{I}$ and GAR class token $\textbf{X}_{gar}$. Prompts, such as bounding boxes and skeletal keypoints, are transformed to prompt tokens $\textbf{F}_{p}$ by the prompt encoder. These tokens, along with instance IDs, are then fed into the recognition decoder for cross-modal updates to get group activity prediction $\textbf{Y}_{gar}$.
Requirement for Full Prompts. Current high-performance group activity recognizers [15, 29, 19] rely on full annotations to leverage all available modalities. This approach, however, makes their performance highly sensitive to the availability and quality of annotations. It is widely acknowledged that obtaining accurate human annotations is difficult. For instance, MP-GCN [32] uses automated tools like YOLOv8 [23] and BoTSORT [1] in NBA dataset [47] processing. However, the resulting annotations may include inaccuracies, such as false positives in background audiences, player ID switching, and ID reassignment upon reappearance. Despite utilizing state-of-the-art object detectors [56, 41], trackers [40, 49], and pose estimatros [45, 16], the lower quality annotations such as inaccurate bounding boxes, missed detections, heavy occlusions, and redundant boxes from unreliable confidence scores may impact the resulting model accuracies and the real-world applicability for GAR.
Lacking Long-term Actor Consistency. Recent studies [55, 32, 15, 29, 19] prioritize short-term actor motions using optical flows or joint displacements between two neighboring frames. While effective for instantaneous activities like ‘spike’, ‘set’, and ‘pass’ in the Volleyball dataset, these methods struggle with longer activities. For instance, differentiating ‘3p-fail-offensive’ from ’3p-succ’ and ‘3p-fail-defensive’ in the NBA dataset [47] requires recognizing a sequence of events: shooter location, ball trajectory, and subsequent ball possession. This highlights the critical importance of long-term consistency when activity duration and input frame length increase in real-world scenarios.
Under-explored Multi-Group Design. Current GAR methods [55, 32, 15, 29, 19] are limited to videos containing a single group, neglecting the common occurrence of multiple groups in real-world scenarios. This limitation, together with the difficulty of acquiring accurate annotations, calls for models capable of area-based prompts to easily select a particular group. Specifically, given a region of interest (ROI), the model should output the group activity label within that area. This becomes even more challenging when groups move dynamically, with only the initial-frame ROI provided. Therefore, developing GAR models capable of accurately processing multi-group scenarios is crucial for bridging the gap between current research and real-world applicability.
@@
Considering the above challenges and motivations, we present PromptGAR, a transformer-based group activity recognition framework that leverages multi-modal prompts (i.e., bounding boxes, skeletons, areas, or their combinations) to achieve high group activity recognition accuracy and input flexibility. (a) Flexible Prompts. It adapts to varying prompt availability, namely full prompts, partial prompts, and no prompts. When comprehensive annotations are present, the full prompt inputs would maximize performance. If only simpler annotations like bounding boxes and instance IDs are available, our model still gives reliable results. Even with no annotations, it can also provide reasonable results using only the raw video. This adaptability makes it suitable for diverse real-world scenarios where annotation quality varies. (b) Flexible Frames. The method has temporal flexibility in accepting videos of varying lengths and frame rates. Furthermore, it effectively recognizes group activities regardless of their temporal location within the video, accommodating both instantaneous actions and sequential events. (c) Flexible Instances. It is designed to handle varying numbers of actors within a scene. Therefore in an ideal scenario, our model can leverage detailed annotations of all individuals for enhanced accuracy. On the otherhand, it can also maintain robust performance even when there are missing annotations for certain actors. (d) Without Retraining. After training with full prompts, PromptGAR automatically gains the flexibility for input prompts, frames, and instances during inference. This significantly enhances its practicality and broadens its applicability to diverse scenarios.
@@
Our implementation is inspired by the Segment Anything [25] by unifying bounding boxes, skeletons, and areas as point prompts and employing a two-way decoder for cross-updating class and prompt tokens. Firstly, the input flexibility is implemented through several design features: MViTv2’s relative positional embedding for variable length of RGB frames, depth-wise prompt pooling to accommodate flexible temporal and modality dimensions, and a robust head that maintains classification stability even when no prompts are available during inference. Secondly, to effectively acquire long-term actor consistency, we introduce relative instance attention, which directly encodes instance IDs. The encoding ensures that the output will remain invariant to instance ID transitions. Lastly, PromptGAR introduces the use of area prompts, enabling the selective recognition of the particular group activity within videos that contain multiple concurrent groups.
@@
Based on the above technical contributions, we evaluate our model under Volleyball [21] and NBA [47] datasets, PromptGAR produces competitive results compared to state-of-the-art GAR methods, showcasing its strong recognition capabilities. Additionally, we demonstrated its flexibility to handle varying prompt inputs, frame sampling, and instance counts while maintaining robust performance without retraining. To validate the ability to recognize specific groups within multi-group scenarios, we also performed experiment in multi-GAR with area prompts by synthetically creating a multi-group dataset. Our main contributions are as follows.
@@
(b) A relative instance attention module for encoding instance IDs and ensuring long-term actor consistency.
(c) Demonstration of capability in multi-group activity recognition using area prompts.
@@ 2.1 Group Activity Recognition
GAR [8] has attracted attention due to its massive success in a variety of real-world applications. Earlier techniques relied heavily on handcrafted features [8, 26, 18, 6, 9, 7, 34] or AND-OR graphs [2, 3, 39]. Recently, methods based on neural network architectures have been widely studied because of their ability to effectively extract features and fuse various modalities.
Various Modalities for GAR. (a) RGB frames and bounding boxes. Existing work either directly crops people from scenes [46] or applies ROIAlign [20] to represent actors from extracted feature maps [50, 38, 29]. (b) Skeletons. Several works [43, 55, 13, 32] explore using human skeleton joints as inputs to avoid substantial computational resources and discrepancies in background and camera settings in video-based approaches. (c) Optical flows. Neighboring RGB frames can produce optical flow images [52]. These flows are then concatenated with RGB frames to create dense inputs to introduce pixel-wise motion information across the temporal dimension [38, 29, 19]. A similar idea is also applied in Composer [55] for skeletal modality by computing the temporal difference of coordinates in two consecutive frames. MP-GCN [32] calculated joint motions and bone motions as extra sparse inputs. (d) Balls. Group labels in the NBA dataset [47], such as ‘2p-succ’ and ‘3p-succ’, are closely related to ball positions; therefore, Composer [55] and MP-GCN [32] treat balls as additional inputs. Howoever, it is difficult to leverage these specific model designs for general scenarios such as non-sports activity recognition. (e) Tracked instance IDs. Recent approaches [50, 29] use optical flows to describe only short-term motions without explicitly encoding instance IDs to preserve long-term actor consistency. We introduce a relative instance ID encoding mechanism to handle this long-neglected modality.
@@
Figure 3: Recognition Decoder. The decoder processes (a) RGB features $\textbf{F}_{I}$ with positional embeddings and (b) the GAR class token $\textbf{X}_{gar}$ and prompt tokens $\textbf{F}_{p}$. (c) The Two-Way Transformer performs cross-updating between these features. (d) Relative instance self-attention, using instance IDs, ensures long-term actor consistency. (e) The GAR head takes updated GAR class token $\widehat{\textbf{X}}_{gar}$ and average of updated prompt tokens $\widehat{\textbf{F}}_{p}$ to predict the group activity label $\textbf{Y}_{gar}$.
2.2 Multi-Group Activity Recognition
Previous research [55, 32, 15, 29, 19] in GAR are limited by having only a single group throughout the entire video. In such a case, all of the modalities, including bounding boxes, skeletons, optical flows, and instance IDs, are only served as ‘enhancers’ to add more clue on determining the group activities. In real-world situations, we are particularly interested in how well recognition works when multiple groups are present in the same video. Similar topics, such as spatio-temporal action detection [17, 28, 30], emphasize on accuracy of temporal intervals and spatial coordinate predictions to recognize individual actions rather than group ones. Therefore, we firstly introduce multi-group activity recognition (Multi-GAR), in which modalities act as both ‘enhancers’ and ‘selectors’ to choose the particular group activity in a video that contains multiple groups. Most GAR methods [55, 32, 15, 29, 19] could be extended to Multi-GAR through bounding box or skeleton offsets. However, those methods are unable to process area prompts, a box defines the region of interest for a group activity. It acts only as a ‘selector’ for a group activity without requiring strong annotations for individual actors. In this paper, we explores multi-GAR problem with the proposed PromptGAR to effectively handles both regular modalities and area prompts.
@@ 3 Method
As shown in Fig. 1, PromptGAR is an end-to-end multi-modal prompt-based framework for group activity recognition. It takes a sequence of frames and corresponding prompts as inputs and outputs the group activity label. In the following subsection, we first introduce how to encode videos and multi-modal prompts in §$\S$3.1. Then we illustrate how to fuse sparse-dense modalities and spatial-temporal information in §$\S$3.2. Finally, we introduce multi-group activity recognition and area prompts in §$\S$3.3.
3.1 Multi-Modality Inputs Encoding
@@
where $\textbf{v}:=[x,y,t]^{T}\in[0,1)^{3}$ is the spatial-temporal coordinate, and $\textbf{B}\in\mathbb{R}^{\lfloor D/2\rfloor\times 3}$ is sampled from a Gaussian distribution $\mathcal{N}(0,1)$. The same positional encoding is also applied to RGB features $\textbf{F}_{I}$. Secondly, we employ learned embeddings for each prompt type $p$, similar to [25]. Thirdly, ID is encoded by relative embeddings, with details in §$\S$ 3.2.
@@
Figure 4: The Relative Instance Attention mechanism that incorporating a constant scale $\epsilon$ and relative instance embeddings $\mathcal{R}_{\text{inst}(i),\text{inst}(j)}$ in the attention block.
The pooling mechanism reduces the computational complexity and avoids out-of-memory (OOM) in the recognition decoder. The depth-wise mechanism, namely the mean operator $\mathcal{M}$ that reduces channels from $T\times 20$ to $1$, is specifically designed for input flexibility across temporal and modality dimensions. Consequently, a model trained on $T_{1}$ frames can perform inference on $T_{2}$ frames $(T_{2}\neq T_{1})$ without requiring architectural modifications or weight changes. Similarly, the model is able to infer with either bounding box prompts alone or skeleton prompts alone, even when trained on both modalities. Note that the instance dimension $N$ is not pooled and is required by ID encoding in §$\S$3.2.
3.2 Cross-Modality Decoding
Leveraging the image embedding $\textbf{F}_{I}$, the prompt features $\textbf{F}_{p}$ and associated instance IDs, PromptGAR takes the GAR class token $\textbf{X}_{gar}$ to decode the GAR logits $\textbf{Y}_{gar}$.
Two-Way Transformer. For simplicity, we refer to these embeddings (not including the image embedding) collectively as ”tokens”.
@@ 3.2 Cross-Modality Decoding
[…]
@@
Relative Instance Attention. Existing research has demonstrated potential in modeling short-term temporal consistency via either optical flows [38, 29, 19] or joint motions [55, 32]. As these methods primarily focus on capturing motion between immediate time steps, the long-term consistency in the movement of objects or individuals is not well handled. Also, the use of absolute instance ID encodings to ensure consistency violates the principle of shift invariance [27]. Namely, the interaction between two prompts becomes dependent on their arbitrary instance IDs, even if they refer to the same underlying object or entity. Inspired by relative positional embedding, we introduce the relative instance ID embedding to address this issue by focusing on whether two tokens belong to the same instance or not.
@@
Here, the number of token pairs belonging to different instances significantly outnumbers those belonging to the same instance. To address the imbalance problem, we assign a learnable embedding $R$ to pairs of tokens within the same instance, while setting the embedding to 00 for pairs from different instances.
$\{R_{\text{inst}(i),\text{ inst}(j)}\}_{i,j}=\left\{\begin{aligned} 0,\quad\text{inst}(i)&\neq\text{ inst}(j)\\ R,\quad\text{inst}(i)&=\text{ inst}(j)\\ \end{aligned}\right.$
Due to the sparsity of $E^{(rel)}$, the scaling factor $\epsilon$ is a constant determined by experiments, with details in §$\S$4.5. Note that relative instance embeddings cannot be applied to cross-attention between image embeddings and tokens, as image embeddings lack associated instance ID information.
@@
where $\text{Mean}(\widehat{\textbf{F}}_{p})\in\mathbb{R}^{D}$ means taking average over $T\times O$.
Figure 5: Left is the 2-Mosaic Volleyball Dataset, showing two vertically concatenated clips. Right is the 4-Mosaic Volleyball Dataset in the $2\times 2$ format. In both cases, a target clip is randomly chosen from each position, and only the annotations corresponding to that target clip are present.
Our approach is also carefully designed to ensure input flexibility, even when the prompts present during the training time are absent during the inference time. During backpropagation, the gradients of $\widehat{\textbf{X}}_{gar}$ and $\text{Mean}(\widehat{\textbf{F}}_{p})$ are identical, indicating they are optimized in the same direction and towards the same representation. Consequently, a well-trained model can function effectively using either $\widehat{\textbf{X}}_{gar}$ or $\text{Mean}(\widehat{\textbf{F}}_{p})$ independently.
3.3 Multi-Group Activity Recognition
We extend the PromptGAR architecture for multi-group activity recognition. Specifically, an area prompt is defined to indicate the region of interest for a particular group activity within a video containing multiple groups. Similar to bounding boxes, it is formulated by three points: the upper-left, center, and lower-right. The prompt encoder then outputs $\textbf{F}_{a}$, a tensor of shape $D\times 1\times O$. The dimension $N=1$ shows that only one area instance is being processed. The recognition decoder, following the same procedure, processes image embeddings $\textbf{F}_{I}$ and concatenated tokens $[\textbf{X}_{gar};\textbf{F}_{a}]$ to generate the prediction $\textbf{Y}_{gar}$. With the spatial awareness provided by ‘Area prompts’, our PromptGAR accurately identifies the target group activity.

### v2

@@ Abstract
We present PromptGAR, a novel framework for Group Activity Recognition (GAR) that offering both input flexibility and high recognition accuracy. The existing approaches suffer from limited real-world applicability due to their reliance on full prompt annotations, fixed number of frames and instances, and the lack of actor consistency. To bridge the gap, we proposed PromptGAR, which is the first GAR model to provide input flexibility across prompts, frames, and instances without the need for retraining. We leverage diverse visual prompts—like bounding boxes, skeletal keypoints, and instance identities—by unifying them as point prompts. A recognition decoder then cross-updates class and prompt tokens for enhanced performance. To ensure actor consistency for extended activity durations, we also introduce a relative instance attention mechanism that directly encodes instance identities. Comprehensive evaluations demonstrate that PromptGAR achieves competitive performances both on full prompts and partial prompt inputs, establishing its effectiveness on input flexibility and generalization ability for real-world applications.
@@ 1 Introduction
Group Activity Recognition (GAR) [7] is fundamentally important for video and event understanding, and it is widely used in areas such as video analytics, human computer interaction, and security systems. GAR processes videos and annotations including bounding boxes, skeletons, ball trajectories, and optical flows to determine the group activity label. Building on prior research [51, 26, 13, 24, 16], we aim to not only enhance group activity recognition accuracy but also to create a more flexible architecture handling diverse prompt, frame, and instance inputs.
@@
Figure 1: PromptGAR Architecture. A sequence of frames II is processed by the video encoder, yielding RGB features $\textbf{F}_{I}$ and GAR class token $\textbf{X}_{gar}$. Prompts, such as bounding boxes and skeletal keypoints, are transformed to prompt tokens $\textbf{F}_{p}$ by the prompt encoder. These tokens, along with instance identities, are then fed into the recognition decoder to get group activity prediction $\textbf{Y}_{gar}$.
Requirement for Full Prompts. Current group activity recognizers [13, 24, 16] rely on full annotations to achieve strong performance at test time. It is widely acknowledged that obtaining accurate annotations is difficult. Even though using state-of-the-art object detectors [52, 35], trackers [34, 44], and pose estimators [40, 14], the lower quality annotations still exist, such as missed detections, redundant boxes with low confidence scores, player ID switching, player ID reassignment upon reappearance, and so on. In real-world scenarios, manually correcting these annotations at test time is often impractical. Instead of being forced to use potentially inaccurate prompts during test time, users in real-world scenarios want the flexibility to choose from full, partial, or no prompts. However, current GAR architectures [51, 26, 13, 24, 16, 45, 30, 32] either do not support partial prompts, or they demand retraining to perform reasonably. So we design our model for the input flexibility: it is trained with full prompts like prior methods, but it delivers competitive results with full prompts and still performs quite well with fewer or no prompts at test time, without retraining.
Requirement for Fixed Frames and Instances. Most current GAR methods [13, 24, 16] are limited to a fixed number of frames and instances as input. Those rigid requirements lead to significant performance degradation in real-world scenarios: (1) As seen in Fig. 2-a, offense or defense relies on who gets the ball at the end. Such key moment can be anywhere in the video, so previous fixed-frame models may miss it depending on how the frames are sampled from a long sequence. (2) Similarly, in Fig. 2-b and c, real-world data often contains missing actors in annotations or false positive detections of spectators as players. Prior fixed-instances models requires a fixed input shape, causing runtime errors or requiring arbitrary padding/truncation that degrades performance. In contrast, our model is designed to accept a flexible number of frames and instances, and it achieves reliable performance without requiring retraining.
Figure 2: Necessity of flexible frames: (a) offense or defense relies on which team gets the ball at the end, and similar key moments can be anywhere in the video. Necessity of flexible instances: (b) false-positives and (c) false-negatives make player counts unfixed.
Lacking Actor Consistency. Recent GAR approaches [49, 26, 51] rely on a fixed player order. Their performance degrades when this order changes. Even though player order in the testing set of either Volleyball [18] or NBA [42] is not supposed to be known, that order is still fixed for both validation and testing in these methods. Consequently, among all epochs, they pick the checkpoint with the best performance under this specific player order. However, in real-world scenarios, such player order is entirely unknown. A robust GAR model, therefore, should give identical performance regardless of the input player order. In our work, we address this by introducing a relative instance attention mechanism that only encodes whether two players are the same or not. This design ensures that player order does not affect our model’s performance at all.
@@
Considering the above challenges and motivations, we present PromptGAR, a transformer-based group activity recognition framework that leverages diverse visual prompts (i.e., bounding boxes, skeletal keypoints, instance identities) to achieve high group activity recognition accuracy and input flexibility. (a) Flexible Prompts. It adapts to varying prompt availability, namely full prompts, partial prompts, and no prompts. When comprehensive annotations are present, the full prompt inputs would maximize performance. If only simpler annotations like bounding boxes and instance identities are available, our model still gives reliable results. Even with no annotations, it can also provide reasonable results using only the raw video. This adaptability makes it suitable for diverse real-world scenarios where annotation quality varies. (b) Flexible Frames. The method has temporal flexibility in accepting videos of varying lengths and frame rates. Furthermore, it effectively recognizes group activities regardless of their temporal location within the video, accommodating both instantaneous actions and sequential events. (c) Flexible Instances. It is designed to handle varying numbers of actors within a scene. Therefore in an ideal scenario, our model can leverage detailed annotations of all individuals for enhanced accuracy. On the other hand, it can also maintain robust performance even when there are missing annotations for certain actors. (d) Without Retraining. After training with full prompts, PromptGAR automatically gains the flexibility for input prompts, frames, and instances during inference. This significantly enhances its practicality and broadens its applicability to diverse scenarios.
@@
Our implementation is inspired by the Segment Anything [21]. We unify bounding boxes and skeletal keypoints as point prompts and employing a two-way decoder for cross-updating class and prompt tokens. Firstly, the input flexibility is implemented through several design features: MViTv2’s relative positional embedding for variable length of RGB frames, depth-wise prompt pooling to accommodate flexible temporal and prompt dimensions, and a robust head that maintains classification stability even when no prompts are available during inference. Secondly, to effectively acquire actor consistency, we introduce relative instance attention, which directly encodes instance IDs. The encoding ensures that the output will remain invariant to instance ID transitions.
@@
Based on the above technical contributions, we evaluate our model under Volleyball [18] and NBA [42] datasets, PromptGAR produces competitive results compared to state-of-the-art GAR methods, showcasing its strong recognition capabilities. Additionally, we demonstrated its flexibility to handle varying prompt inputs, frame sampling, and instance counts while maintaining robust performance without retraining. Our main contributions are as follows.
@@
(b) A relative instance attention module for encoding instance identities and ensuring actor consistency.
@@ 2.1 Group Activity Recognition
GAR [7] has attracted attention due to its massive success in a variety of real-world applications. Earlier techniques relied heavily on handcrafted features [7, 22, 15, 5, 8, 6, 28] or AND-OR graphs [2, 1, 33]. Recently, methods based on neural network architectures have been widely studied because of their ability to effectively extract features and fuse various visual prompts.
Various Prompts for GAR. (a) RGB frames and bounding boxes. Existing work either directly crops people from scenes [41] or applies ROIAlign [17] to represent actors from extracted feature maps [45, 32, 24]. (b) Skeletons. Several works [37, 51, 11, 26] explore using human skeleton joints as inputs to avoid substantial computational resources and discrepancies in background and camera settings in video-based approaches. (c) Optical flows. Neighboring RGB frames can produce optical flow images [47]. These flows are then concatenated with RGB frames to create dense inputs to introduce pixel-wise motion information across the temporal dimension [32, 24, 16]. A similar idea is also applied in Composer [51] for skeletal information by computing the temporal difference of coordinates in two consecutive frames. MP-GCN [26] calculated joint motions and bone motions as extra sparse inputs. (d) Balls. Group labels in the NBA dataset [42], such as ‘2p-succ’ and ‘3p-succ’, are closely related to ball positions; therefore, Composer [51] and MP-GCN [26] treat balls as additional inputs. However, it is difficult to leverage these specific model designs for general scenarios such as non-sports activity recognition. (e) Tracked instance identities. Recent approaches [45, 24] use optical flows to describe only short-term motions without explicitly encoding instance identities to preserve actor consistency. We introduce a relative instance identities encoding mechanism to handle this long-neglected prompt.
Figure 4: Recognition Decoder. The decoder processes (a) RGB features $\textbf{F}_{I}$ with positional embeddings and (b) the GAR class token $\textbf{X}_{gar}$ and prompt tokens $\textbf{F}_{p}$. (c) The Two-Way Transformer performs cross-updating between these features. (d) Relative instance self-attention, using instance identities, ensures actor consistency. (e) The GAR head takes updated GAR class token $\widehat{\textbf{X}}_{gar}$ and average of updated prompt tokens $\widehat{\textbf{F}}_{p}$ to predict the group activity label $\textbf{Y}_{gar}$.
@@ 3 Method
As shown in Fig. 1, PromptGAR is an end-to-end prompt-based framework for group activity recognition. It takes a sequence of frames and corresponding prompts as inputs and outputs the group activity label. In the following subsection, we first introduce how to encode videos and visual prompts in $\S$3.1. Then we illustrate how to fuse various visual prompts and spatial-temporal information in $\S$3.2.
3.1 Visual Inputs Encoding
@@
where $\textbf{v}:=[x,y,t]^{T}\in[0,1]^{3}$ is the spatial-temporal coordinate, and $\textbf{B}\in\mathbb{R}^{\lfloor D/2\rfloor\times 3}$ is sampled from a Gaussian distribution $\mathcal{N}(0,1)$. The same positional encoding is also applied to RGB features $\textbf{F}_{I}$. Secondly, we employ learned embeddings for each prompt type $p$, similar to [21]. Thirdly, ID is encoded by relative embeddings, with details in $\S$ 3.2.
@@
The pooling mechanism reduces the computational complexity and avoids out-of-memory (OOM) in the recognition decoder. The depth-wise mechanism, namely the mean operator $\mathcal{M}$ that reduces channels from $T\times 20$ to $1$, is specifically designed for input flexibility across temporal and prompt dimensions. Consequently, a model trained on $T_{1}$ frames can perform inference on $T_{2}$ frames $(T_{2}\neq T_{1})$ without requiring architectural modifications or weight changes. Similarly, the model is able to infer with either bounding box prompts alone or skeleton prompts alone, even when trained on both prompts. Note that the instance dimension $N$ is not pooled and is required by relative identities encoding in $\S$3.2.
3.2 Cross-Visual Decoding
Leveraging the image embedding $\textbf{F}_{I}$, the prompt features $\textbf{F}_{p}$ and associated instance identities, PromptGAR takes the GAR class token $\textbf{X}_{gar}$ to decode the GAR logits $\textbf{Y}_{gar}$.
Figure 5: The Relative Instance Attention mechanism that incorporating a constant scale $\epsilon$ and relative instance identity embeddings $\mathcal{R}_{\text{inst}(i),\text{inst}(j)}$ in the attention block.
Two-Way Transformer. For simplicity, we refer to these embeddings (not including the image embedding) collectively as “tokens”.
@@ 3.2 Cross-Visual Decoding
$\widehat{\textbf{F}}_{I},[\widehat{\textbf{X}}_{gar};\widehat{\textbf{F}}_{p}]$ $=\text{Two-Way}(\texttt{images},\texttt{tokens})$
@@
Relative Instance Attention. Existing research has demonstrated potential in modeling short-term temporal consistency via either optical flows [32, 24, 16] or joint motions [51, 26]. As these methods primarily focus on capturing motion between immediate time steps, the long-term consistency in the movement of objects or individuals is not well handled. Also, the use of absolute instance ID encoding to ensure consistency violates the principle of shift invariance [23]. Namely, the interaction between two prompts becomes dependent on their arbitrary instance IDs, even if they refer to the same underlying object or entity. Inspired by relative positional embedding, we introduce the relative instance identity embedding to address this issue by focusing on whether two tokens belong to the same instance.
@@
Here, the number of token pairs belonging to different instances significantly outnumbers those belonging to the same instance. To address the imbalance problem, we assign a learnable embedding $R$ to pairs of tokens within the same instance, while setting the embedding to 0 for pairs from different instances.
[…]
Due to the sparsity of $E^{(rel)}$, the scaling factor $\epsilon$ is a constant determined by experiments, with details in $\S$4.5. Note that relative instance embeddings cannot be applied to cross-attention between image embeddings and tokens, as image embeddings lack associated instance ID information.
@@
where $\text{Mean}(\widehat{\textbf{F}}_{p})\in\mathbb{R}^{D}$ means taking average over $T\times O$. Our approach is also carefully designed to ensure input flexibility, even when the prompts present during the training time are absent during the inference time. During backpropagation, the gradients of $\widehat{\textbf{X}}_{gar}$ and $\text{Mean}(\widehat{\textbf{F}}_{p})$ are identical, indicating they are optimized in the same direction and towards the same representation. Consequently, a well-trained model can function effectively using either $\widehat{\textbf{X}}_{gar}$ or $\text{Mean}(\widehat{\textbf{F}}_{p})$ independently.